compare and contrast quadratic and exponential function (using visuals/diagrams) . In your discussion, consider using the following talking points: first differences, quadratic, exponential, equation, graph, f(x)=x^2, f(x)=2^x. 



1
Expert's answer
2021-02-04T14:34:03-0500

Firstly, quadratic function is a polynomial with the highest degree of 2. That is, "x^2" is always present there. While, exponential function has a variable as it exponent e.g. "2^x".


Differences:

When the independent variable changes by a constant amount, quadratic functions have constant second differences, while exponential functions have a constant ratio.


Graph:

Taking "f(x)=x^2" and "f(x)=2^x" to be the example for quadratic and exponential function respectively. We see the graph as follows

This is the graph of "f(x)=x^2"



This is the graph of "f(x)=2^x"



We see that the quadratic graph makes/form a Parabola (whether facing up or down). While the exponential function graph rises or falls freely in One Direction.